OpenStudy (anonymous):

A ball is dropped from a height of 300 feet. Its velocity after t seconds is v=-32t ft/s A) How fast is the ball dropping after 4 seconds (i got -128, i want to make sure) B)Determine the position function C) How far has the ball dropped after 4 seconds D)How many seconds will it take for the ball to hit the ground

OpenStudy (shayaan_mustafa):

lol. it classical physics. you should post in physics group.

OpenStudy (anonymous):

...this is calculus

OpenStudy (shayaan_mustafa):

is this? which book?

OpenStudy (shayaan_mustafa):

ok let us solve this. be here.

OpenStudy (anonymous):

calculus seventh dition

OpenStudy (shayaan_mustafa):

howard anton?

OpenStudy (anonymous):

Larson Hostetler Edwards

OpenStudy (amistre64):

the velocity is slowing down after you drop it?

OpenStudy (amistre64):

acceleration due to gravity is defined in feetric measure as: -32ft/^2 a constant rate so: a(t) = -32 velocity is the integration of acceleration: {S} a(t) = v(t) = -32t + c , and we are given that at t=4 v= -32 -32 = -32(4) + c -32+32(4) = c 32(4-1) = c = 96 v(t) = -32t +96 the integration of velocity gives us position {S} v(t) = h(t) = -16t^2 +96t +c

OpenStudy (anonymous):

no the negative would mean that the ball is heading toward the earth

OpenStudy (amistre64):

yeah, i saw that afterwards :) when time (t) = 0 we are at a height of 300 sooo 300 = -16(0)^2 + 96(0) + c ; hence c = 300 h(t) = -16t^2 +96t +300

OpenStudy (amistre64):

i got some of the information mixed up fromthis tiny little screen

OpenStudy (amistre64):

the 96 is bad; spose we are simply dropping the ball; giving it no initial velocity; the that part goes to 0

OpenStudy (amistre64):

h(t) = -16t^2 + 300 v(t) = -32t a(t) = -32

OpenStudy (amistre64):

so v(4) = -128

OpenStudy (amistre64):

the rest of it is just filling in the numbers i believe

OpenStudy (anonymous):

so what is the velocity in the position function?

OpenStudy (amistre64):

initial velocity is zero inthe position function

OpenStudy (amistre64):

if we throw the ball; that gives it an inital velocity other than zero; simply letting it drop imparts no velocity and lets gravity do all the work

OpenStudy (anonymous):

ok i got it thanks

OpenStudy (amistre64):

yw